A random variable X has the cumulative distribution function given as
F (x) =
8><>:
0; for x < 1
x2 − 2x + 2
2 ; for 1 ≤ x < 2
1; for x ≥ 2
Calculate the variance of X
1
Expert's answer
2020-05-19T20:16:55-0400

P(1<x<2) = F(2) - F(1)= (22 -2*(2) + 2)/2 - (12 - 2*(1) + 2)/2= (4-4+2)/2 - (1-2+2)/2 = 2/2 - 1/2= 1/2

So if 1/2 of the pdf lies between 1<x<=2, none of it is x>2, and none of it is below x<1, the remaining 1/2 of the pdf must lie at x=1, so

f(x)=F'(x)={1/2,x=1;x-1,1<x<2;0 - otherwise

E(x)=1/2+∫21x(x-1)dx=4/3= 1.333

E(x2)=1/2+∫21x2(x-1)dx=23/12=1.916

VAR=|E(x2)-E(x)2|=0.139